If a $200 pair of shoes is reduced by 15%, what is the discount?
Found 2 solutions by lynnlo, ikleyn:
Answer by lynnlo(4176)
Answer by ikleyn(53765)

.
If a $200 pair of shoes is reduced by 15%, what is the discount?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~


Then the discount is   0.15*200 = 15*2 = 30  dollars.         ANSWER